# Bayesian Model Parameter Learning in Linear Inverse Problems: Application in EEG Focal Source Imaging

### v1

@@ 5.5.2 Skull Conductivity Estimates
5.5.3 Reference Method: Alternating Dipole Scan
As a reference, we compared our dipole source and skull conductivity estimates with the corresponding estimates of the following alternating approach
$\begin{split}\hat{x}_{l}^{(k+1)}&\leftarrow\min_{i=1:n}\{\min_{x_{i}}\|L_{e}(v-A(\sigma^{(k)})^{i}x_{i}-e_{*})\|_{2}^{2}\},\\ \hat{\sigma}^{(k+1)}&\leftarrow\min_{\sigma}\|L_{e}(v-A(\sigma)^{l}\hat{x}_{l}^{(k+1)}-e_{*})\|_{2}^{2},\end{split}$ (37)
where $A(\sigma)^{l}\approx A(\sigma^{(k)})^{l}+\frac{\partial A^{l}}{\partial\sigma}|_{\sigma=\sigma^{(k)}}(\sigma-\sigma^{(k)})$.
6 Results and Discussion
@@
Random white noise was added to the measurements. In Section 6.1, a low noise level, $\mathrm{SNR}=40$ dB, was used in order to study the effects of the varying source amplitude on the results. In Section 6.2, a higher noise level, $\mathrm{SNR}=30$ dB, was used.
All the dipole source reconstructions were carried out using the standard leadfield model (in the coarse mesh) that had fixed skull conductivity $\sigma_{0}=0.0103$ S/m; the alternating dipole scan obviously updated the leadfield model during the iterations and used the standard leadfield model only in the initial step. For the model parameter (skull conductivity) estimation, CG+Iter also updated the leadfield model in the iterations.
@@
6.1 Effects of Varying Source Amplitude in Source Localization Improvement and Skull Conductivity Estimates
In the first test case, we simulated boundary measurements resulting from 88 radial dipoles (one at a time) with increasing source amplitude (from 0.3 to 4.2, in normalized source amplitude units), and then computed source reconstructions and skull conductivity estimates using the proposed approaches, CG (35), CG+Iter. (Algorithm 1), and GP (36).
Figure 4 shows the dipole localization errors, $X_{\mathrm{st}}$ and $X_{\mathrm{BAE}}$, of the standard solution (33) and the BAE solution (34), respectively, for four different source intensities. As can be seen, the localization errors of the BAE solutions are much more often smaller than the localization errors of the standard solution in all the cases. The benefits of BAE are more evident in the case with $\sigma_{\mathrm{true}}=0.00601\,\mathrm{S/m}<\sigma_{0}$ than in the case with $\sigma_{\mathrm{true}}=0.0139\,\mathrm{S/m}>\sigma_{0}$. Moreover, the localization errors do not seem to have any particularly strong trend with respect to the source intensity.
Figure 4: Localization errors of the standard solution and the BAE solution in the two studied cases, $\sigma_{\mathrm{true}}=0.00601$ S/m and $\sigma_{\mathrm{true}}=0.0139$ S/m, when different source intensities were used. The standard model used in all the reconstructions assumed (erroneously) skull conductivity $\sigma_{0}=0.0103$ S/m.
@@
In Figure 5, we show with box plots the source localization improvements ($\Delta X=X_{\mathrm{st}}-X_{\mathrm{BAE}}$, in milli meters) and the estimated skull conductivities in the two test cases as a function of the source amplitude. The medians ar represented by the horizontal lines inside the boxes, the edges of the boxes denote the 25th and 75th percentiles, and the whiskers and outliers indicate the range of values. It is worth emphasizing that the source reconstructions were carried out by first, using the standard leadfield model that assumed (erroneously) skull conductivity value $\sigma_{0}=0.0103$ S/m, and second, using the same standard model accompanied with the statistics of the BAE modeling. Higher $\Delta X$ values in the first (left) image of Fig. 5 indicate greater improvements using the proposed BAE dipole scan. Light gray refers to the $\sigma_{\mathrm{true}}<\sigma_{0}$ case and dark gray to the $\sigma_{\mathrm{true}}>\sigma_{0}$ case. We can see that localization improvements can be achieved using the proposed BAE dipole scan in both cases, when compared to using the standard dipole scan, and that the BAE dipole scan particularly improves the results in the $\sigma_{\mathrm{true}}<\sigma_{0}$ case.
The second (middle) and third (right) image in Fig. 5 show box plots of the estimated skull conductivities in the test cases with $\sigma_{\mathrm{true}}=0.00601\,\mathrm{S/m}<\sigma_{0}$ and $\sigma_{\mathrm{true}}=0.0139\,\mathrm{S/m}>\sigma_{0}$, respectively. The skull conductivities were estimated using the proposed techniques, CG, CG+Iter and GP, across various source amplitudes. The horizontal solid blue line represents the standard conductivity $\sigma_{0}$, and the dashed black line shows the true conductivity used in the test cases.
Based on these results, the CG-based conductivity estimation performs well only when the source amplitude is close to the mode of the source amplitude distribution $\pi(|x_{i}|)$ (see Appendix A1 for an explanation). Furthermore, the solution tends to be close to the standard value ($\sigma_{0}$) when the source amplitude is small. This is because for small source amplitudes, the reconstructed error coefficient ($\alpha_{i}$) is usually small (due to the scaling effect, see Section 4.2) and thus, favors conductivity values close to $\sigma_{0}$ (as was expected based on the red curve in Fig. 2). Furthermore, we can see that the skull conductivity estimates for increasing dipole amplitude either decrease (in the $\sigma_{\mathrm{true}}<\sigma_{0}$ case), or increase (in the $\sigma_{\mathrm{true}}>\sigma_{0}$ case). Therefore, it is clear that the conductivity solutions of CG depend on the amplitude of the source.
@@
6.2 Source Localization Improvements and Skull Conductivity Estimates with the Proposed Algorithms
@@
In the following two case studies, we fixed the source amplitude to 1.3 (normalized units). We estimated the sources and their localization accuracy, as before. For the model parameter (skull conductivity) estimation, we used only the proposed CG+Iter and GP, as it was shown in the previous section that these two were superior to the simple CG approach. We ran the algorithms for each location with 5 different (measurement) noise realizations at SNR = 30 dB, and we estimated the corresponding localization improvements, $\Delta X=X_{\mathrm{st}}-X_{\mathrm{BAE}}$, and the errors of the conductivity estimates as percentages, $|\Delta\sigma|=100\times\frac{|\hat{\sigma}-\sigma_{\mathrm{true}}|}{\sigma_{\mathrm{true}}}$. The presented results are the average values of $\Delta X$ and $|\Delta\sigma|$ for each tested location. As a reference method, we computed the source reconstructions and skull conductivity estimates by using the alternating dipole scan (37).
@@
Fig. 6 depicts the results of the proposed Bayesian approach for the $\sigma_{\mathrm{true}}<\sigma_{0}$ case applied over selected source locations in MRI slices 72, 74, 76, and 78. The top row presents the source localization improvements resulting from the BAE solution with respect to the standard model. The white triangles depict the locations in which the source localization improvements were the highest ($\Delta X>6$ mm), and the red and orange triangles correspond to smaller localization improvements, ranging from 2 mm to 6 mm. In locations marked with white squares, the localization errors were negligibly small ($\leq 2$ mm). The yellow (upside down) triangles depict negative localization improvements (between --6 mm and --2 mm), which means that in these cases the standard model (with erroneous skull conductivity value) performed better than the proposed BAE approach. The note below the top row gives the percentage of cases in which the localization improvements were greater than 2 mm.
@@
Figure 6: Results of the proposed Bayesian approaches in terms of source localization improvement and conductivity estimation error across MRI slices (72, 74, 76, and 78) when $\sigma_{\mathrm{true}}<\sigma_{0}$ and SNR = 30 dB. The top row displays localization improvements, with triangles indicating different magnitudes of the improvements (in milli meters). As the conductivity values were estimated together with the sources, both the localization improvements and conductivity estimation errors are presented with the help of the (true) locations of the sources that generated the corresponding EEG testing data. The middle and bottom rows show the errors of the conductivity estimates when CG+Iter. (middle) and GP (bottom) were used. The symbols represent different levels of conductivity errors (in percentages): white squares denote errors less than 5%, and circles with shades of blue and black higher skull conductivity estimation errors. Here, GP achieved conductivity errors below 5% in 71.6% of cases, outperforming CG+Iter. that achieved the same only in 48.9% of cases.
@@
In Fig. 7, we show the results of the reference method, the alternating dipole scan set in (37), in the $\sigma_{\mathrm{true}}<\sigma_{0}$ case. The localization improvements and conductivity estimation errors are presented using the same notations and symbols as in Fig. 6. Figs. 8 and 9 present the corresponding Bayesian and reference results in the $\sigma_{\mathrm{true}}>\sigma_{0}$ case.
@@ 6.3.1 Source Estimates
The Bayesian approximation error approach showed substantial improvements in source localization accuracy when compared to both, using the standard model with fixed skull conductivity and the alternating dipole scan that was used as a reference method. When compared to the standard model, the BAE improved source localization by $\geq 2$ mm in 75% of the tests in the $\sigma_{\mathrm{true}}<\sigma_{0}$ case and in 51% of the tests in the $\sigma_{\mathrm{true}}>\sigma_{0}$ case. In the rest of the tests, the performances of the BAE and the standard model were similar, with the exception of only two source locations in which the standard model outperformed the BAE. The alternating dipole scan performed in most tests equally to the standard model, outperforming the standard model only in 37% of the tests in the $\sigma_{\mathrm{true}}<\sigma_{0}$ case and in 19% of the tests in the $\sigma_{\mathrm{true}}>\sigma_{0}$ case.
The improvements provided by the BAE modeling were particularly clear in the case where the true skull conductivity was lower than $\sigma_{0}$. This is because the EEG measurements, and thus the approximation errors, are non-linearly proportional to the skull conductivity, roughly $\varepsilon\propto 1/\sigma$. Therefore, the absolute values of the approximation errors in the $\sigma_{\mathrm{true}}<\sigma_{0}$ case were (on average) higher than in the $\sigma_{\mathrm{true}}>\sigma_{0}$ case. In the tests where the approximation errors were high (causing poor performance for the standard model), the BAE approach could effectively alleviate the source localization errors with the help of the embedded model-based training statistics of the approximation errors. The other way around, when the approximation errors were smaller in the $\sigma_{\mathrm{true}}>\sigma_{0}$ case, also the potential improvements in the localization accuracy that could be gained with the help of the BAE approach were more modest. These findings are in line with our previously published results [31].
@@ 6.3.2 Model Parameter Estimates
GP offered an effective alternative to the above as it did not require iterations. However, the prerequisite was that a suitable kernel function was needed to describe the relationship between the model parameter and $\varepsilon$. In our case, this relationship could be approximated based on the underlying physics (see Appendix A2). From the tested approaches, GP gave the best model parameter estimates.
@@
6.4 Model-based Learning in Inverse Problems

### v2

@@ 5.1.8 Skull Conductivity Estimates
Reference Method: Alternating Single Dipole Scanning:
We compared our estimates of dipole source and skull conductivity with the estimates of a reference method based on alternating minimization [14, 15]. The corresponding estimates were the following
$\begin{split}\hat{x}_{l}^{(k+1)}&\leftarrow\min_{i=1:n}\{\min_{x_{i}}\|L_{e}(v-A(\sigma^{(k)})^{i}x_{i})\|_{2}^{2}\},\\ \hat{\sigma}^{(k+1)}&\leftarrow\min_{\sigma}\|L_{e}(v-A(\sigma)^{l}\hat{x}_{l}^{(k+1)})\|_{2}^{2},\end{split}$ (37)
where $A(\sigma)^{l}\approx A(\sigma^{(k)})^{l}+\frac{\partial A^{l}}{\partial\sigma}|_{\sigma=\sigma^{(k)}}(\sigma-\sigma^{(k)})$ where the derivative was evaluated as in [62].
5.2 Simulated Experiments and Results
@@
Random white noise was added to the measurements. In Section 5.2.1, a low noise level, $\mathrm{SNR}=40$ dB, was used in order to study the effects of the varying source amplitude on the results. In Section 5.2.2, a higher noise level, $\mathrm{SNR}=30$ dB, was used.
All the dipole source reconstructions were carried out using the standard leadfield model (in the coarse mesh) that had fixed skull conductivity $\sigma_{0}=0.0103$ S/m; the alternating dipole scanning obviously updated the leadfield model during the iterations and used the standard leadfield model only in the initial step. For the model parameter (skull conductivity) estimation, CG+Iter also updated the leadfield model in the iterations.
@@
5.2.1 Effects of Varying Source Amplitude in Source Localization Improvement and Skull Conductivity Estimates
In the first test case, we simulated boundary measurements stemming from 88 radial dipoles (one at a time) with increasing source amplitude (from 0.3 to 4.2, in normalized source amplitude units [^5] Instead of using the typical single dipole source amplitude values, 10-50 nAm, we used the normalized ones for numerical stability.), and then computed source reconstructions, using (33) and (34), and skull conductivity estimates using the proposed approaches, CG (35), CG+Iter. (Algorithm 1), and GP (36).
Figure 4 shows the histograms of dipole localization errors, $X_{\mathrm{st}}$ and $X_{\mathrm{BAE}}$, of the standard solution (33) and the BAE solution (34), respectively, for the four different source intensities. Each histogram line corresponds to results of 88 simulated sources with a fixed source intensity.
As can be seen, the localization errors of the BAE solutions are much more often smaller than the localization errors of the standard solution in all the cases. Moreover, the localization errors do not seem to have any particularly strong trend with respect to the source intensity.
Figure 4: Histogram lines of the localization errors of the standard solution and the BAE solution in the two studied cases, $\sigma_{\mathrm{true}}=0.00601$ S/m and $\sigma_{\mathrm{true}}=0.0139$ S/m, when different source intensities were used. The standard model used in all the reconstructions assumed (erroneously) skull conductivity $\sigma_{0}=0.0103$ S/m.
@@
In Figure 5, we show with box plots the source localization improvements ($\Delta X=X_{\mathrm{st}}-X_{\mathrm{BAE}}$, in milli meters) and the estimated skull conductivities in the two test cases as a function of the source amplitude. The medians are represented by the horizontal lines inside the boxes, the edges of the boxes denote the 25th and 75th percentiles, and the whiskers and outliers indicate the range of values. It is worth emphasizing that the source reconstructions were carried out by first, using the standard leadfield model that assumed (erroneously) skull conductivity value $\sigma_{0}=0.0103$ S/m, and second, using the same standard model accompanied with the statistics of the BAE modeling. Higher $\Delta X$ values in the first (left) image of Fig. 5 indicate greater improvements using the proposed BAE dipole scanning. Light gray refers to the $\sigma_{\mathrm{true}}<\sigma_{0}$ case and dark gray to the $\sigma_{\mathrm{true}}>\sigma_{0}$ case. We can see that localization improvements can be achieved using the proposed BAE dipole scanning in both cases, when compared to using the standard dipole scanning. The benefits of BAE are more evident in the case with $\sigma_{\mathrm{true}}=0.00601\,\mathrm{S/m}<\sigma_{0}$ than in the case with $\sigma_{\mathrm{true}}=0.0139\,\mathrm{S/m}>\sigma_{0}$.
The second (middle) and third (right) image in Fig. 5 show box plots of the estimated skull conductivities in the test cases with $\sigma_{\mathrm{true}}=0.00601\,\mathrm{S/m}<\sigma_{0}$ and $\sigma_{\mathrm{true}}=0.0139\,\mathrm{S/m}>\sigma_{0}$, respectively. The skull conductivities were estimated using the proposed techniques, CG, CG+Iter and GP, across various source amplitudes. The horizontal dashed black line represents the standard conductivity $\sigma_{0}$, and the solid blue line shows the true conductivity used in the test cases.
Based on these results, the CG-based conductivity estimation performs well only when the source amplitude is close to the mode of the source amplitude distribution $\pi(|x_{i}|)$ (see Appendix A2 for an explanation). Furthermore, the solution tends to be close to the standard value ($\sigma_{0}$) when the source amplitude is small. This is because for small source amplitudes, the reconstructed error coefficient ($\alpha_{i}$) is usually small (due to the scaling effect, see Section 4.2) and thus, favors conductivity values close to $\sigma_{0}$ (as was expected based on the red curve in Fig. 2). Furthermore, we can see that the skull conductivity estimates of CG for increasing dipole amplitude either decrease (in the $\sigma_{\mathrm{true}}<\sigma_{0}$ case), or increase (in the $\sigma_{\mathrm{true}}>\sigma_{0}$ case). Therefore, we observe here the interaction effect (explained in Sections 3.1 and 3.3.1) based on which a change in the source amplitude can hinder the conductivity estimation with CG, even though, based on the underlying physics, the source amplitude does not depend on the skull conductivity.
@@
5.2.2 Source Localization Improvements and Skull Conductivity Estimates with the Proposed Algorithms
@@
In the following two case studies, we fixed the source amplitude to 1.3 (normalized units). We estimated the sources and their localization accuracy, as before. For the model parameter (skull conductivity) estimation, we used only the proposed CG+Iter and GP, as it was shown in the previous section that these two were superior to the simple CG approach. We ran the algorithms for each location with 5 different (measurement) noise realizations at SNR = 30 dB, and we estimated the corresponding localization improvements, $\Delta X=X_{\mathrm{st}}-X_{\mathrm{BAE}}$, and the errors of the conductivity estimates as percentages, $|\Delta\sigma|=100\times\frac{|\hat{\sigma}-\sigma_{\mathrm{true}}|}{\sigma_{\mathrm{true}}}$. The presented results are the average values of $\Delta X$ and $|\Delta\sigma|$ for each tested location. As a reference method, we computed the source reconstructions and skull conductivity estimates by using the alternating dipole scanning (37).
@@
Fig. 6 depicts the results of the proposed Bayesian approach for the $\sigma_{\mathrm{true}}<\sigma_{0}$ case applied over selected source locations in MRI slices 72, 74, 76, and 78. The top row presents the source localization improvements resulting from the BAE solution with respect to the standard model. The white triangles depict the locations in which the source localization improvements were the highest ($\Delta X>6$ mm), and the red and orange triangles correspond to smaller localization improvements, ranging from 2 mm to 6 mm. In locations marked with white squares, the localization errors were negligibly small ($\leq 2$ mm) discretization considered. The yellow (upside down) triangles depict negative localization improvements (between --6 mm and --2 mm), which means that in these cases the standard model (with erroneous skull conductivity value) performed better than the proposed BAE approach. The note below the top row gives the percentage of cases in which the localization improvements were greater than 2 mm.
@@
Figure 6: Results of the proposed Bayesian approaches in terms of source localization improvement and conductivity estimation error across coronal MRI slices (72, 74, 76, and 78) when $\sigma_{\mathrm{true}}<\sigma_{0}$ and SNR = 30 dB. The top row displays localization improvements, with triangles indicating different magnitudes of the improvements (in milli meters). As the conductivity values were estimated together with the sources, both the localization improvements and conductivity estimation errors are presented with the help of the (true) locations of the sources that generated the corresponding EEG testing data. The middle and bottom rows show the errors of the conductivity estimates when CG+Iter. (middle) and GP (bottom) were used. The symbols represent different levels of conductivity errors (in percentages): white squares denote errors less than 5%, and circles with shades of blue and black higher skull conductivity estimation errors. Here, GP achieved conductivity errors below 5% in 61.4% of cases, outperforming CG+Iter. that achieved the same only in 40.9% of cases.
@@
In Fig. 7, we show the results of the reference method, the alternating dipole scanning set in (37), in the $\sigma_{\mathrm{true}}<\sigma_{0}$ case. The localization improvements and conductivity estimation errors are presented using the same notations and symbols as in Fig. 6. Figs. 8 and 9 present the corresponding Bayesian and reference results in the $\sigma_{\mathrm{true}}>\sigma_{0}$ case.
@@ 5.3.1 Source Estimates
The Bayesian approximation error approach showed substantial improvements in source localization accuracy when compared to both, using the standard model with fixed skull conductivity and the alternating dipole scanning that was used as a reference method. When compared to the standard model, the BAE improved source localization by $\geq 2$ mm in 78.4% of the tests in the $\sigma_{\mathrm{true}}<\sigma_{0}$ case and in 51.1% of the tests in the $\sigma_{\mathrm{true}}>\sigma_{0}$ case. In the rest of the tests, the performances of the BAE and the standard model were similar; there were only two source locations in which the standard model outperformed the BAE. The alternating dipole scanning performed in most tests equally to the standard model, outperforming the standard model only in 37.5% of the tests in the $\sigma_{\mathrm{true}}<\sigma_{0}$ case and in 19.3% of the tests in the $\sigma_{\mathrm{true}}>\sigma_{0}$ case.
The improvements provided by the BAE modeling were particularly clear in the case where the true skull conductivity was lower than $\sigma_{0}$. This is because the EEG measurements, and thus the approximation errors, are non-linearly proportional to the skull conductivity, roughly $\varepsilon\propto 1/\sigma$. Therefore, the absolute values of the approximation errors in the $\sigma_{\mathrm{true}}<\sigma_{0}$ case were (on average) higher than in the $\sigma_{\mathrm{true}}>\sigma_{0}$ case. In the tests where the approximation errors were high (causing poor performance for the standard model), the BAE approach could effectively alleviate the source localization errors with the help of the embedded model-based training statistics of the approximation errors. The other way around, when the approximation errors were smaller in the $\sigma_{\mathrm{true}}>\sigma_{0}$ case, the potential improvements in the localization accuracy that could be gained with the help of the BAE approach were more modest. These findings are in line with our previously published results [31].
@@ 5.3.2 Model Parameter Estimates
GP offered an effective alternative to the above as it did not require iterations. However, the prerequisite was that suitable mean and kernel functions were needed to describe the relationship between the model parameter and $\varepsilon$. In our case, this relationship could be approximated based on the underlying physics (see Appendix A3). From the tested approaches, GP gave the best model parameter estimates.
@@
5.3.3 General Remarks, Potential and Limitations
In this work, by applying the dipole scanning approach that assumes strict single-dipole activity, we successfully removed the dependence of the approximation error on the source location. This was achieved by computing location-specific error estimates; in other words, the error depended solely on the dipole source and the skull conductivity at a fixed location. As a result, the modeling error effects could be described using a single eigenvector, and the computational effort and complexity of the estimation process were reduced considerably in the inversion.
Furthermore, the single-dipole scanning approach employed in our demonstrations has been shown to be equivalent to sLORETA which is a well-established and widely used technique for localizing single smooth source activities [60]. This result enables direct application of the approach to scenarios involving smoother, spatially extended single sources. In particular, visualizing the sLORETA estimate would be equivalent to visualizing a goodness-of-fit distribution defined by $1-\frac{\|v-A_{0}^{(i)}\hat{x}^{i}-w_{i}\hat{\alpha}_{i}-\varepsilon_{i*}\|_{2}}{\|v\|_{2}}$ for $i=1\ldots,n$, where $(\hat{x}_{i},\hat{\alpha_{i}})$ are the estimates obtained from the single-dipole scan.
We note that the design of prior distributions (e.g. Gaussian, Laplace or hierarchal priors) in the BAE posterior (10) of the EEG source imaging problem is not a straightforward task. This due to the fact that the EEG source imaging problem suffers from an inherent limitation, namely the so-called depth-bias [68] which means that the reconstructed active sources typically locate close to the brain surface, regardless whether the true source activity is on the brain surface or deeper in the brain. Unfortunately, this depth-bias problem has not yet been fully resolved; one usually tries to overcome it by using either depth weights and/or hyperparameters in the prior distribution [69, 70, 71, 72, 73, 74, 75, 76, 77, 78]. In this work, by considering strictly single-dipole activity, we eliminated the need for these weighting schemes and specially designed priors.
Here, we want to acknowledge that low signal-to-noise ratios can impact both primary unknowns and error estimates, and consequently the accuracy of model parameter inference. However, the specific tolerance to noise depends on properties of the inverse problem, such as the number of measurements and the characteristics of the convolution kernel. In this study, we noticed that with SNR = 30 dB we could still recover meaningful reconstructions. In our previous study [31], we observed that the benefits of BAE were negligible when SNR = 20 dB. As detailed in Appendix A4, one way to test this would be to compare the noise level of the measurements with the magnitude of the expected approximation errors [42].
One practical observation we made in the current work was that it is preferable to use an average model parameter value (skull conductivity) in the standard model, since it facilitates model parameter inference under Gaussian assumptions and faster convergence when linearized model parameter estimates are later performed. If the model parameter estimation step is not needed, then the choice of the standard model may not always be very critical [54].
Finally, as the BAE approach without model parameter estimation has been used in various (linear and non-linear) problems [45, 46, 47, 48, 49, 50, 51, 52, 53, 54], we would expect that in many similar cases as these the joint inference of the primary unknown and error approximation (11) would be tractable. Subsequently, based on the available training data, the properties (physics) of the problem at hand, and the complexity of the approximation error term, different inference strategies for the model parameters could be formulated. In addition to the CG, and the proposed CG+Iter. and GP, others such as polynomial chaos expansions and equivalent techniques could be tested for the model parameter inference.
6 Conclusion and Future Work